OpenStudy (anonymous):

A planet has a mass of about 6.4 x 1023 kg, and its moon has a mass of about 9.6 x 1015 kg. If the magnitude of the gravitational force between the two bodies is 4.6 x 1015 N, how far apart are the planet and its moon? a.9.4 X 106 c.4.6 X 109 b.2.6 X 1015 d.1.2 X 102

OpenStudy (anonymous):

\(\LARGE F_g=\frac{Gm_1m_2}{r^2}\) ---> \(\LARGE r=\sqrt {\frac{Gm_1m_2}{F_g}}\) substitute all your given values. Welcome to Openstudy!